OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 4) and (−2, 13)? 3x + y = 7 3x − y = 7 x − 3y = 7 x + 3y = 7

Parth (parthkohli):

I can help here.

OpenStudy (anonymous):

Okay, thanks. :)

Parth (parthkohli):

The points are always given as (x,y). Do you know that?

OpenStudy (anonymous):

Yes.

Parth (parthkohli):

This says that when x is 1, then y is 4. We can check right?

OpenStudy (anonymous):

Right.

Parth (parthkohli):

Let's start checking with the first equation and see if it is satisfied.

OpenStudy (anonymous):

The first one is right.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow3(1) + 4 = 7 }\) \(\Large \color{MidnightBlue}{\Rightarrow 3 + 4 = 7 }\) \(\Large \color{MidnightBlue}{\Rightarrow 7 = 7 }\) Stage one is complete.

Parth (parthkohli):

Yes, it is correct :D

OpenStudy (anonymous):

Thanks for talking me through that. :) Think you could help me with another one?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 3(-2) + 13 = 7 }\) \(\Large \color{MidnightBlue}{\Rightarrow -6 + 13 = 7 }\) \(\Large \color{MidnightBlue}{\Rightarrow 7 = 7 }\) \(\Huge \checkmark\)

Parth (parthkohli):

And sure!

OpenStudy (anonymous):

Awesome. :) A company builds computers. It costs $7,600 to build 10 computers and $12,200 to build 20 computers. Which equation models the cost, C(x), as a linear function of the number of computers built, x? C(x) = 460x − 3000 C(x) = 3000x − 460 C(x) = 460x + 3000 C(x) = 3000x + 460

Parth (parthkohli):

You can again check here.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

How would I go about that?

Parth (parthkohli):

I see that the third one is correct.

OpenStudy (anonymous):

Would I fill in the cost for C(x)?

OpenStudy (anonymous):

Thank you. :)

Parth (parthkohli):

Well, you see that 10 computers cost 7600, so you can see --> 460 * 10 + 3000 = 7600

OpenStudy (anonymous):

Thanks so much! ♥